OpenStudy (anonymous):

3x-4y=-1 and x+y=5 substitution..

OpenStudy (anonymous):

x = 5 - y 3x - 4y = -1 3( 5 - y ) -4y = -1

OpenStudy (anonymous):

try to continue the process.